You have $100 invest for 1 year at a rate of 10%. Bread sells for $2 today and that means you can buy 50 loaves today. Bread is expected to have a 4% inflation next year. What is your real rate of return? How many loaves of bread can you buy next year?
Answer #1

Real rate of return=(1+10%)/(1+4%)-1=5.77%

Loaves of bread to be bought next year=100/2*(1+5.77%)=52.8846 or 52